OpenStudy (anonymous):

How can one sixthx − 5 = one fifthx + 2 be set up as a system of equations? 6y − x = −30 5y − x = 10 5y + x = −30 5y + x = 10 6y − 6x = −30 5y − 5x = 10 5y + 5x = −30 5y + 5x = 10 @Directrix plz help me

OpenStudy (princeharryyy):

put x = -210

OpenStudy (princeharryyy):

take like terms on one side and solve for x @iamabarbiegirl

OpenStudy (anonymous):

would it be A?

OpenStudy (princeharryyy):

yes it is @ @iamabarbiegirl

OpenStudy (princeharryyy):

A

OpenStudy (princeharryyy):

correct ! @iamabarbiegirl

OpenStudy (anonymous):

yeah 1/6 x -5=1/5x+2

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

this is the correct problem

OpenStudy (princeharryyy):

@iamabarbiegirl take equation 1 that is this 1/6 x -5 and put it equal to y, we get... 1/6 x -5 =y which is equal to 6y − x = −30 take equation 2 that is this 1/5x+2 and put it equal to y, we get... 1/5x+2 =y which is equal to 5y − x = 10 and hence the answer is A... they have just equated the equation with y... when u equal y from both equations ... u get 1/6 x -5=1/5x+2

OpenStudy (princeharryyy):

and ofcourse Yes th question is correct... thats why I agreed with your answer @iamabarbiegirl

OpenStudy (anonymous):

yeah im just confused now haha i just wanna really make sure its correct

Directrix (directrix):

y = (1/6) x - 5 and y = (1/5) x + 2 6y = x - 30 and 5y = x + 10

Directrix (directrix):

Those are the same as this: 6y − x = −30, 5y − x = 10 which is set A.

OpenStudy (anonymous):

yayyyy so my answer was right?!?!

Directrix (directrix):

Yes, A. Sorry it took me so long to get with the program.

OpenStudy (anonymous):

np